OpenStudy (anonymous):

I need help on about 3 questions about parallel and perpendicular lines and equations. Somebody please help?

OpenStudy (anonymous):

Question 1: Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). (2 points) y = 1/3 x + 7 y = -3x + 3 y =1/3 x + 3 y = -3x + 7

OpenStudy (anonymous):

@amistre64 @Ashleyisakitty @allie_bear22 @Broganleab @Charlie_xD @canadalover_ Can anybody help please?!!!!

OpenStudy (anonymous):

HAHAHAHAHHAHAHA U THINK THAT YOU ARE SO STRONG I BET

OpenStudy (anonymous):

YOU HAVE NO PECS THIS IS GR8

OpenStudy (anonymous):

I need help with Algebra. This has nothing to do with working out. I don't need you opinions and no, I don't think I'm strong. I have muscles that I work hard to get. You are just jealous.

OpenStudy (anonymous):

NICE HAT TOO!!

OpenStudy (anonymous):

SWOLL????????

OpenStudy (anonymous):

GO CHEAT ON YOUR ONLINE SCHOOL SOMEWHERE ELSE

OpenStudy (anonymous):

Can you just stop? I need somebody to explain to me how to do this. I'm not going to explain myself to you. Either help me or get off of my question. Thanks.

OpenStudy (anonymous):

YOUR CHEST IS SO SMALLLLL

OpenStudy (anonymous):

Omg, I don't care about your opinions! Help me or leave! It's not a hard thing to do.

OpenStudy (anonymous):

Alright, so how do we tell how a line is parallel? Do you know that?

OpenStudy (anonymous):

GOOODDDDBYEEEE

OpenStudy (anonymous):

Well, when you graph both of the lines they will be next to each other in a way and don't they have the same slope?

OpenStudy (anonymous):

Exactly, they have the same slope. The equation they gave you was in Slope-intercept form, so we don't even have to convert it! First thing we have to do is create a new equation with an equal sope, which will be: \[y =-3x - b\]

OpenStudy (anonymous):

Now you have to solve for b, by entering in a point.

OpenStudy (anonymous):

So, \[6 = -3 \times -1 + b\] I'm sure you can do that on your own.

OpenStudy (anonymous):

Oh, okay thank you!

OpenStudy (anonymous):

I can solve the other questions too now :)

OpenStudy (anonymous):

Awesome, sorry about that typo with the - b, that's wrong.